What is the volume of the largest rectangular parallelepiped which can be inscribed in the ellipsoid x2/9+y2/16+z2/36=1? Show that the answer you get gives you the largest volume.(DO NOT USE LAGRANGE MULTIPLIERS)


1
Expert's answer
2021-12-14T13:48:39-0500

let (x,y,z) be the coordinates of the corner of the box that is inscribed in an ellipsoid. Thus "x,y,z \geq0" and equation of ellipsoid is

"\frac{x^2}{9}+\frac{y^2}{16}+\frac{z^2}{36}=1"


"\frac{z^2}{36}=1-\frac{x^2}{9}-\frac{y^2}{16}"

so "z=6\sqrt{1-\frac{x^2}{9}-\frac{y^2}{16}}"


The volume of the box is represented as

"V=(2x)(2y)(2z)=48xy\sqrt{1-\frac{x^2}{9}-\frac{y^2}{16}}"

For "x\geq0,y\geq0, and \ (\frac{x^2}{a^2})+(\frac{y^2}{b^2})\leq1"


For analysis it is easier to deal with "V^2\ than\ V" as it becomes more complex to find the partial derivative in terms of x and y

"V^2=2304[x^2y^2-\frac{x^4y^2}{9}-\frac{x^2y^4}{16})"


Since "V=0 \ if \ x=0 \ or \ y=0 \ or\ (\frac{x^2}{a^2})+(\frac{y^2}{b^2})=1,"

the maximum value of V2 and hence of V will occur at critical point of V2 to find the critical point we we have to find the partial derivative in terms of x and y


"\frac{\delta V^2}{\delta x}=2304[2xy^2-\frac{4x^3y^2}{9}-\frac{2x^2y^4}{16})"


"=4608xy^2(1-\frac{2x^2}{9}-\frac{y^2}{16})........(1)"


similarly


"\frac{\delta V^2}{\delta y}=4608xy^2(1-\frac{x^2}{9}-\frac{2y^2}{16})........(2)"


Now we will equate the above partial derivatives to zero so as to find critical points


"(1-\frac{2x^2}{9}-\frac{y^2}{16})=0\\\frac{2x^2}{9}+\frac{y^2}{16}=1........(3)"


"(1-\frac{x^2}{9}-\frac{2y^2}{16})=0\\\frac{x^2}{9}+\frac{2y^2}{16}=1........(4)"


Now by equating (3) and (4) we get


"\frac{2x^2}{9}+\frac{y^2}{16}=1=\frac{x^2}{9}+\frac{2y^2}{16}"


which

"\frac{x^2}{9}=\frac{1}{3}\implies x=\frac{9}{\sqrt{3}}"


"\frac{y^2}{16}=\frac{1}{3}\implies y=\frac{16}{\sqrt{3}}"


The volume of the largest box that can be inscribed inside the ellipsoid is


"V=48xy\sqrt{1-\frac{x^2}{9}-\frac{y^2}{16}}"


"=48(\frac{9}{\sqrt{3}})(\frac{16}{\sqrt{3}})\sqrt{1-\frac{1}{3}-\frac{1}{3}}"


"=1330.215"